OpenStudy (anonymous):

Calculate the length of the side x, given that tan θ = 0.4 will post bellow

OpenStudy (anonymous):



OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

I don't know the formula for tangent

OpenStudy (igreen):

SOH CAH TOA Tangent = Opposite/Adjacent

OpenStudy (igreen):

So in this case it will be: \(\sf tan(0.4) = \dfrac{x}{15}\)

OpenStudy (anonymous):

6cm?

OpenStudy (igreen):

Yep, that's if you round.. First we find tan(0.4), which is 0.422793219: \(\sf 0.422793219 = \dfrac{x}{15}\) Now we multiply 15 to both sides: \(\sf x = 0.422793219 \times 15\) \(x = 6.34189829 \approx 6\)

OpenStudy (anonymous):

thank you very much. again :)